8. The reserve requirement, open market operations, and the moneysupply Assume that banks do not hold excess reserves and thaWhich of the following statements help to explain why, in the real world, the Fed cannot precisely control the money supply?
Answer #1

Total reserves of the banking system are equal to 300. There is no currency or excess reserves. Simple deposit multiplier = 1
Answer #2

(1)

Money multplier (m) = 1 / Reserve requirement (r)

Money supply = Deposit x m

(i) When r = 5% = 0.05

m = 1 / 0.05 = 20

Money supply = $300 x 20 = $6,000

(ii) When r = 10% = 0.1

m = 1 / 0.1 = 10

Money supply = $300 x 10 = $3,000

(2) Higher reserve requirement is associated with a Smaller money supply.

(3) When Fed wants to increase money supply by $200 and r = 10% (i.e. m = 10),

Fed will use open market operations to Buy $20 worth (= $200 / m = $200 / 10 = $20) government bonds.

(4) When r increases from 10% to 25%,

Increase in r causes money multiplier to fall to 4 (= 1 / 0.25). Under these conditions, Fed will have to Buy $50 worth (= $200 / m = $200 / 4 = $50) government bonds.

(5) Reasons why Fed cannot precisely control the money supply are:

- Fed cannot control whether and to what extent banks hold excess reserves (The more (less) excess reserves banks hold back, the lower (higher) the money supply).

- Fed cannot control the amount of money households choose to hold as currency (The higher (lower) the currency drainage, the lower (higher) the MM, and the lower (higher) the money supply)


answered by: Allen